### v2

# M4FC: a Multimodal, Multilingual, Multicultural, Multitask Real-World Fact-Checking Dataset

Abstract
Existing real-world datasets for multimodal fact-checking have multiple limitations: they contain few instances, focus on only one or two languages and tasks, suffer from evidence leakage, or rely on external sets of news articles for sourcing true claims. To address these shortcomings, we introduce M4FC, a new real-world dataset comprising 4,982 images paired with 6,980 claims. The images, verified by professional fact-checkers from 22 organizations, represent a diverse range of cultural and geographic contexts. Each claim is available in one or two out of ten languages. M4FC spans six multimodal fact-checking tasks: visual claim extraction, claimant intent prediction, fake image detection, image contextualization, location verification, and verdict prediction. We provide baseline results for all tasks and analyze how combining intermediate tasks influences verdict prediction performance. We make our dataset and code available.111github.com/UKPLab/M4FC
M4FC: a Multimodal, Multilingual, Multicultural, Multitask
Real-World Fact-Checking Dataset
Jiahui Gengthanks: These authors contributed equally to this work.1, Jonathan Tonglet*2,3,4, Iryna Gurevych1,2 1 Mohamed bin Zayed University of Artificial Intelligence (MBZUAI), UAE 2Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science, TU Darmstadt and National Research Center for Applied Cybersecurity ATHENE, Germany 3 Department of Electrical Engineering, KU Leuven, Belgium 4 Department of Computer Science, KU Leuven, Belgium www.ukp.tu-darmstadt.de
1 Introduction
Since 2020, approximately 80% of online misinformation verified by fact-checkers has been multimodal, combining text with images, videos, or audio (Dufour et al., 2024). Debunking multimodal misinformation is time-consuming for human fact-checkers (Silverman, 2013; Khan et al., 2024). It typically involves performing several sub-tasks and utilizing specialized tools, like satellite and street-view imagery (Mossou and Higgins, 2021; Khan et al., 2023, 2024). Additionally, it requires expert knowledge of specific languages and cultures (Silverman, 2013). This work targets the two most common categories of image-based misinformation (Dufour et al., 2024): (1) out-of-context (OOC) misinformation, where authentic images are misrepresented by false claims paired; and (2) misinformation using manipulated or fake images.
Figure 1: An out-of-context instance from M4FC. The six AFC tasks are shown in blue and their ground truth output in grey. Purple boxes indicate (partially) incorrect LLM answers, while green boxes indicate correct ones.
Multimodal automated fact-checking (AFC) assists human fact-checkers by automating parts of their workflow (Akhtar et al., 2023). Several datasets have been proposed for multimodal AFC. Synthetic datasets, on the one hand, are large and scalable (Luo et al., 2021; Shao et al., 2024) but cannot fully mimic real-world misinformation. On the other hand, real-world datasets consist of claims verified by human fact-checkers, but these datasets tend to be much smaller (Zlatkova et al., 2019; Hu et al., 2023; Papadopoulos et al., 2024; Tonglet et al., 2024). Existing real-world datasets suffer from additional limitations. While multimodal misinformation is a global problem, these datasets cover only one or two languages and lack geographic diversity. Furthermore, the human fact-checking workflow involves several intermediate tasks before providing a verdict. However, existing datasets fail to account for most of these tasks.
In this work, we introduce M4FC, a multimodal, multilingual, multicultural, and multitask real-world fact-checking dataset, with 4,982 images and 6,980 claims. M4FC addresses several limitations of existing resources. It contains labels for six multimodal AFC tasks organized into a pipeline, as illustrated in Figure 1. This pipeline allows for automating a larger part of the fact-checking workflow and studying the interaction between intermediate tasks and verdict prediction. Images, claims, task labels, and metadata are drawn from articles produced by 22 fact-checking organizations across 17 countries. As a result, M4FC offers broad cultural and geographic coverage and spans 10 claim languages: Arabic, Dutch, English, French, German, Portuguese, Spanish, Tamil, Telugu, and Turkish. We introduce two new multimodal AFC tasks. Visual claim extraction addresses the mandatory step of formulating verifiable claims based on a screenshot of a social media post containing the image. Location verification validates candidate locations for images using two types of evidence previously overlooked in multimodal AFC research: maps and satellite images. We provide baseline results for all tasks. Furthermore, we evaluate how different sequential combinations of intermediate tasks affect verdict prediction performance.
In summary, our contributions are as follows: (1) we introduce M4FC, a large-scale real-world multimodal, multilingual, multicultural, and multitask fact-checking dataset; (2) we introduce two new challenging multimodal AFC tasks; and (3) we conduct comprehensive experiments, including analyses of how combining intermediate tasks impacts verdict prediction performance.
2 Related work
Claims Evidence
Dataset # tasks # images # claims # languages Data split Source Leak? Type
Weibo (Jin et al., 2017) 1 9,528 9,528 1 (Zh) Event FC & News - -
Fauxtography (Zlatkova et al., 2019) 1 1,233 1,233 1 (En) Random FC & News No Text
COSMOS (test) (Aneja et al., 2023) 1 1,700 3,400 1 (En) No split FC & News - -
MR2 (Hu et al., 2023) 1 14,700 14,700 2 (En, Zh) Event FC & News Yes Text, Image
Ookpik (Pham et al., 2024) 1 545 1,090 1 (En) No split FC & News - -
Post-4V (Geng et al., 2024) 1 186 186 1 (En) No split FC - -
VERITE (Papadopoulos et al., 2024) 1 662 1,000 1 (En) No split FC Yes Text, Image
5Pils (Tonglet et al., 2024) 1 1,676 1,676 1 (En) Temporal FC No Text
5Pils-OOC (Tonglet et al., 2025) 2 624 1,248 1 (En) No split FC No Text, Image
AVerImaTeC (Cao et al., 2025) 1 1,297 1,297 1 (En) Temporal FC No Text, Image
XFacta (Xiao et al., 2025) 1 2,400 2,400 1 (En) Random CN & News Unknown Text, Image
M4FC (ours) 6 4,982 6,980 10 Temporal FC No Text, Map, Satellite
Table 1: Comparison of real-world multimodal AFC datasets. En and Zh indicate claims in English and Chinese, while FC, News, and CN stand for fact-checking articles, news articles, and X Community Notes, respectively.
2.1 Synthetic multimodal AFC datasets
Several synthetic datasets have been proposed for multimodal AFC (Sabir et al., 2018; Müller-Budack et al., 2020; Luo et al., 2021; Shao et al., 2024; Liu et al., 2025). While synthetic datasets are large and scalable, they often suffer from biases. Papadopoulos et al. (2025) demonstrated that simple models can exploit these biases to achieve SOTA performance, raising concerns about the reliability of synthetic data for multimodal AFC.
2.2 Real-world multimodal AFC datasets
Real-world datasets, summarized in Table 1, collect claims from fact-checking articles. Because human fact-checking is time-intensive, limited articles are available, and real-world datasets are smaller than synthetic ones. They also often lack true claims (Nielsen and McConville, 2022), as fact-checkers prioritize articles about false claims. To compensate, many datasets sample image-caption pairs from reputable news outlets and treat them as true claims (Jin et al., 2017; Zlatkova et al., 2019; Aneja et al., 2023; Hu et al., 2023). However, this introduces a distribution mismatch (Nielsen and McConville, 2022), as the content, resolution, and topics of news images may differ from those in online misinformation. Existing datasets cover one or two claim languages and provide labels for one task: verdict prediction. Exceptions include 5Pils (Tonglet et al., 2024), which addresses the task of image contextualization, and 5Pils-OOC (Tonglet et al., 2025), which covers both tasks. Many datasets gather evidence via reverse image search (RIS) (Zlatkova et al., 2019; Tonglet et al., 2024) or by querying a search engine (Hu et al., 2023; Papadopoulos et al., 2024; Cao et al., 2025). To ensure realistic evaluation, evidence written by fact-checking organizations or published after the claim date should be excluded, a filtering step applied by some datasets (Zlatkova et al., 2019; Tonglet et al., 2024; Cao et al., 2025). However, this cannot prevent the risk that leaked evidence are part of the pre-training corpus of MLLMs (Xiao et al., 2025).
M4FC addresses several limitations of existing datasets. By gathering data from 22 fact-checking organizations from around the world, it covers 10 languages and is geographically diverse. Furthermore, M4FC includes six tasks, allowing for the combination of intermediate tasks with verdict prediction and aligning multimodal AFC research more closely with human fact-checking practices.
3 M4FC Dataset
3.1 Six multimodal AFC tasks
M4FC includes six tasks, including two new ones, visual claim extraction and location verification.
Claims are not always expressed online in an explicit and verifiable format. Visual claim extraction (VCE) generates an English-language claim from the screenshot of a social media post. This task involves two challenges: (1) extracting visual cues and text from a wide range of languages, and (2) leveraging external knowledge. As shown in Figure 2, parts of the claim appear in the text, others on the building, and further clarification requires external knowledge.
Claimant intent prediction (CIP) seeks to identify the underlying purpose of the claimant, such as expressing sarcasm or promoting a political agenda. We formulate it as a free-text generation task. While Da et al. (2021) and Wu et al. (2025) predicted claimant intents using synthetic data, we extend their work to real-world claims.
Fake image detection (FD) identifies images that are manipulated or fake. Manipulations include face swaps or background changes, while fake images encompass AI-generated content and human-made digital artwork. This task has two labels: authentic and manipulated/fake.
Image contextualization (IC) determines the original context of an image (Tonglet et al., 2024). This task is framed as predicting a set of items. We cover the following context items (Tonglet et al., 2024, 2025): provenance (binary variable indicating whether the image has previously appeared on the web), source (original author or publisher), date, location, motivation (why the source made or published the image), people, things, and event.
Human fact-checkers often compare the image with aerial views of a candidate location to further validate it (Khan et al., 2024). Inspired by this practice, we define the task of location verification (LV): given a map or satellite view of a candidate location, the goal is to determine whether it is consistent with the target image. In the example of Figure 3, the first candidate location, Sart Canal Bridge, is confirmed by the matching road network below the bridge. In contrast, another candidate, the Ringvaart aqueduct, is ruled out because it spans a four-lane highway, which does not match the content of the image. This task relates to cross-view geolocalization (Lin et al., 2013), where the objective is to retrieve a matching satellite image from a large database, given a ground-level image.
Verdict prediction (VP) classifies a claim as true or false. The task is defined in four settings, based on two factors: whether claims are English-only or multilingual, and whether the verdicts are imbalanced or balanced. The multilingual setting covers 10 languages. The imbalanced setting captures the real-world distribution of fact-checking verdicts, in which most claims are false. The balanced setting replaces a portion of false claims with synthetic true claims based on the image’s true context. The English-only imbalanced setting can be used for training, while other settings are only recommended for evaluation, due to the presence of linguistic biases, as discussed in Appendix A.
Figure 2: Illustration of the visual claim extraction task. Blue, yellow, and gray boxes indicate visual cues, external knowledge, and the extracted claim, respectively.
3.2 Dataset construction
3.2.1 Data collection
The data in M4FC is labeled by human fact-checking experts. We collected fact-checking articles from 22 organizations, based in 17 different countries, all of which are members of the International Fact-Checking Network (IFCN), ensuring adherence to professional standards. Table 2 reports the distribution of images per fact-checking organization. Appendix E provides the geographic distribution of the images. M4FC images span six continents and 159 countries, resulting in greater cultural diversity and reduced geographic bias compared to prior work. The top three continents represented are Asia, Africa, and Europe.
All task labels and metadata are provided in the human-written raw article text. Following Tonglet et al. (2024), we utilize GPT4o (OpenAI, 2023) to extract task labels and metadata from the article, if available. Afterwards, GPT4o generates additional synthetic true claims based on the context items of images labeled as false (out-of-context). For a random half of those instances, the original false claim is replaced by the synthetic true claim under the balanced verdict prediction setting. All prompts are provided in Appendix B.
Images suitable for location verification are identified semi-automatically. We use keywords to filter articles referencing maps, satellites, street names, or landmarks. Furthermore, the image needs to have a location context item that can be mapped to coordinates. We then manually verify that the image shows an outdoor scene. Finally, GPT4o generates a list of likely locations for the image, and the first two with valid coordinates serve as negative candidate locations.
Finally, we apply a temporal dataset split to avoid leakage (Glockner et al., 2022). Appendix C provides more details about data collection.
3.2.2 Data validation
Figure 3: Illustration of the location verification task. Satellite images are sourced from ESRI World Imagery.
For each organization, we randomly sample up to 20 articles. Following Tonglet et al. (2024), each article is reviewed by two annotators who evaluate the correctness of the data fields extracted by GPT4o. Annotators are recruited through the Prolific platform. 79 annotators participated in the study, of which 33 were excluded for failing more than five out of 15 attention checks. A label is considered incorrect if both annotators agree on it. The majority of labels are accurate, with scores ranging from 86.2% (for claimed date) to 99.31% (for verdict). We compute inter-annotator agreement (IAA) using Randolph’s $\kappa$ (Randolph, 2005), which ranges from 0.32 to 0.81 across all data fields, indicating moderate to strong agreement. If we exclude instances labeled by GPT4o as not “enough information”, the main source of disagreement, the IAA ranges from 0.45 to 0.85. Guidelines and detailed scores are provided in Appendices D.1 and D.2.
3.2.3 Dataset statistics
Table 3 provides the distribution of instances per task. 1,148 images are unsuitable for visual claim extraction because they do not contain any embedded text to infer the claim. Location verification has the fewest instances because it applies to a specific subset of images that show outdoor scenes. More statistics are provided in Appendix E. We release a high-quality subset with labels validated by all annotators, as explained in Appendix F.
Organization Country Language # articles
factly.in India English, Kannada, Telugu 1027 (20.6%)
fatabyyano.net Jordan Arabic 658 (13.2%)
pesacheck.org Kenya English, French 561 (11.3%)
snopes.com USA English 454 (9.1%)
indiatoday.in India English 333 (6.7%)
leadstories.com USA English 328 (6.6%)
poligrafo.sapo.pt Portugal Portuguese 284 (5.7)%)
congocheck.net DRC French 260 (5.2%)
larepublica.pe Peru Spanish 253 (5.1%)
correctiv.org Germany German 220 (4.4%)
dfrac.org India English 104 (2.1%)
knack.be Belgium Dutch 95 (1.9%)
mythdetector.ge Georgia English 67 (1.3%)
dubawa.org Nigeria English 65 (1.3%)
youturn.in India Tamil 59 (1.2%)
dogrula.org Turkey Turkish 52 (1.0%)
211check.org South Sudan English 51 (1.0%)
verify-sy.com Turkey Arabic 42 (0.8%)
tjekdet.dk Denmark Danish 29 (0.6%)
factcheckcenter.jp Japan Japanese 18 (0.4%)
factchecker.gr Greece Greek 17 (0.3%)
factchecklab.org Hong Kong Chinese 5 (0.1%)
Table 2: M4FC data source distribution.
3.2.4 Evidence collection
For image contextualization and verdict prediction, we collect up to 20 web evidence using the Google RIS engine.222cloud.google.com/vision/docs/detecting-web For location verification, we use the coordinates of the candidate locations to collect maps from OpenStreetMap (OpenStreetMap contributors, 2017) and satellite images from ESRI World Imagery,333arcgis.com/home/item.html accessed with staticmap.444github.com/komoot/staticmap
4 Experiments
4.1 Evaluation metrics
For visual claim extraction and claimant intent prediction, we assess the generated texts using Rouge-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTScore (Zhang et al., 2020). We also adopt G-Eval Liu et al. (2023b), a normalized LLM-based metric aligned with human judgment, to evaluate consistency between predictions and ground truth. We use GPT4o as the LLM. Appendix G provides the G-Eval prompts. For fake image detection, location verification, and verdict prediction, we report the macro F1-score (F1). For image contextualization, following Tonglet et al. (2025), we adopt METEOR for source, motivation, things, and event, F1 for people, $\Delta$ for date, and $CO\Delta$ for location. $\Delta$ and $CO\Delta$ measure the inverse proportional distance from the ground-truth in years and in thousand kilometers, respectively. For verdict prediction with intermediate tasks, we report the accuracy (Acc), precision (Pre), recall (Rec) and F1.
4.2 Baselines
Task # instances # train # dev # test Input
VCE 3,839 2,285 415 1,144 I
CIP 1,789 1,113 164 512 I, C
FD 4,954 2,978 498 1,478 I
IC 4,720 2,887 480 1,353 I, C, E
LV 195 96 27 72 I, E
VP 4,955 2,979 498 1,478 I, C, E
Table 3: Number of instances and input types per task. I is the image, C is the claim, E is the evidence set.
For fake image detection, we evaluate three specialized detectors, Deep-Fake Detector (Prithiv, 2024), AI-Image Detector (Dafilab, 2025), and RealFake Detector (Mikedata, 2025), considering both their pre-trained versions and fine-tuned variants on the MF4C training set. For other tasks, we employ three widely adopted open-weight MLLMs, InternVL2.5-8B (Chen et al., 2024), Qwen2.5-VL-7B (Qwen-Team, 2025), and Llama-3.2-11B-Vision (MetaAI, 2024), alongside two proprietary models, GPT4o-mini (OpenAI, 2023) and Gemini-1.5-Flash (Gemini-Team, 2024), with task-specific prompts provided in Appendix H. In addition, for image contextualization and verdict prediction, we benchmark against the current state-of-the-art (SOTA) methods, COVE (Tonglet et al., 2025) and DEFAME (Braun et al., 2025), respectively. COVE combines two models: LlavaNext-7B (Liu et al., 2023a, 2024) and Llama-3-8B (MetaAI, 2024). For DEFAME, we use GPT4o-mini as the LLM. Appendix I provides implementation details.
4.3 Baselines results
We analyze results for all tasks on the test set. Additional result tables are provided in Appendix J.
(a) Visual claim extraction.
(b) Claimant intent prediction.
(c) Fake image detection.
(d) Verdict prediction.
Figure 4: Visual claim extraction (VCE), claimant intent prediction (CIP), fake image detection (FD), and verdict prediction (VP) results (%). All VP results are shown in the balanced setting.
Figure 5: Visual claim extraction error examples with Gemini-1.5-Flash on M4FC test.
Visual claim extraction
Figure 4(a) shows that most models achieve decent performance, with the exception of Llama-3.2-11B-Vision, while Gemini-1.5-Flash consistently outperforms all others. Rouge-L and METEOR scores show the token-level alignment between generated and reference claims, whereas BERTScore and G-Eval indicate a more moderate level of semantic correspondence. We manually analyzed 50 randomly sampled predictions from Gemini-1.5-Flash. Among them, 38 claims are accurate, nine were partially correct but lacked external world knowledge, and three were incorrect due to the model’s inability to capture sarcasm in social media posts. Figure 5 presents two error cases: on the left, the model fails to connect the image with its broader context, although the claim publication date and the landscape hint at a connection with the Maui fires; on the right, the model overlooks the sarcasm and extracts a claim that directly contradicts the ground truth.
Claimant intent prediction
Figure 4(b) shows that Gemini-1.5-Flash achieves the best performance, with a clear margin on BERTScore and G-Eval compared to other models. Llama-3.2-11B trails slightly behind. GPT4o-mini underperforms because it often refuses to execute the task, flagging claimant intent prediction as a potential policy violation. The overall scores are low, highlighting that real-world claimant intent prediction remains challenging for current MLLMs.
Fake image detection
Figure 4(c) reports F1 scores for specialized fake detectors. While the pre-trained Deep-Fake detector achieves a relatively strong F1 of 70.31%, AI-Image detector is almost ineffective at only 6.41%. Although the M4FC train set is small and does not overlap in time with the test set, fine-tuning on it yields high performance gains across all models. This means that features can be learned from prior real-world cases and generalized to newer cases.
Image contextualization
As shown in Table 4, performance varies substantially across different context items, with no model dominating all categories. Qwen2.5-VL-7B achieves the highest scores on Date (30.3%), Location (45.0%), Motivation (16.8%), and Event (16.2%). In contrast, Gemini-1.5-Flash demonstrates strength in recognizing entities, leading in People (45.0%) and Things (22.7%). The large gaps observed in Location and People further reveal that these dimensions are particularly challenging for current MLLMs. COVE does not provide better results than most MLLMs. On the contrary, it is outperformed by all MLLMs for Source and Date. All models struggle with Source. Consistent with prior work, predicting the date of the image is more challenging than predicting its location (Tonglet et al., 2024, 2025).
Source Date Loc. Mot. People Things Event
[…]
InternVL-2.5-8B 4.8 21.6 26.5 9.5 24.4 13.4 8.0
Qwen2.5-VL-7B 5.2 30.3 45.0 16.8 41.0 13.7 16.2
Llama-3.2-11B 7.7 29.1 38.6 10.5 34.6 14.0 12.2
GPT4o-mini 5.7 20.4 28.9 7.6 20.4 13.3 8.3
Gemini-1.5-Flash 7.1 23.9 42.3 16.6 45.0 22.7 15.8
COVE (Llama-3-8B) 1.7 17.3 33.0 16.2 34.5 12.1 14.4
Table 4: Image contextualization (IC) results (%). Loc., Mot., and M stand for Location, Motivation, and METEOR, respectively.
Location verification
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] & [Uncaptioned image]
[…]
Table 5: Location verification (LV) F1 scores (%). [Uncaptioned image] is the candidate location as a text, while [Uncaptioned image] and [Uncaptioned image] are the corresponding map and satellite views.
We evaluate MLLMs by comparing each claim image with candidate locations presented in three formats: text, map, satellite, or both map and satellite. As shown in Table 5, F1 is often lower with aerial views compared to the text baseline, in particular for commercial models. This suggests that many MLLMs can recognize the location but struggle to match the image to its corresponding aerial view. Model behaviors further diverge by modality: Llama-3.2-11B performs worse than random (49.18%) when given the candidate location as text, while Qwen2.5-VL-7B achieves its best scores with satellite imagery. Gemini-1.5-Flash consistently leads across all settings, and GPT4o-mini is the only model to benefit from combining map and satellite views. Despite these differences, overall F1 scores remain modest, underscoring the difficulty of this multi-image reasoning task. Error cases in Figure 6 illustrate common failure modes: the dam example shows the model’s difficulty in distinguishing fine-grained structural details, while the aqueduct example highlights limitations of using a fixed zoom level (15), as discussed further in Appendix K.
Verdict prediction
Figure 4(d) compares models under both multilingual and English-only balanced settings. Proprietary models show strong linguistic robustness, maintaining comparable performance between multilingual and English claims, regardless of RIS evidence availability. By contrast, open-weight models display much larger disparities across languages. InternVL2.5-8B and Llama-3.2-11B, in particular, perform markedly worse on multilingual claims, with Llama-3.2-11B suffering the steepest drop. Interestingly, Qwen2.5-VL-7B shows the opposite trend, achieving slightly better results in multilingual settings. Across all systems, providing RIS evidence consistently improves results, and the effect is most pronounced for models struggling with multilingual inputs. These results highlight substantial variation in cross-lingual generalization: while leading proprietary models achieve near-linguistic parity, many open-weight alternatives remain challenged in multilingual contexts. Overall performance levels also underscore the difficulty of verdict prediction: no model exceeds 80% F1. The English-only performance of GPT4o-mini with DEFAME or with RIS evidence does not differ much. This suggests that RIS constitutes the most useful source of evidence, and other evidence search tools introduced by the DEFAME framework, like web search, image search, and geolocation, have a limited impact on performance.
Intermediate Tasks GPT4o-mini Qwen2.5-VL-7B
VCE CIP FD IC LV ER Acc Pre Rec F1 Acc Pre Rec F1
60.92 72.41 71.41 71.91 54.84 66.66 71.05 68.79
53.62 73.14 53.40 61.73 49.26 64.52 58.80 61.53
60.92 72.61 70.98 71.78 53.20 66.03 62.73 64.34
61.31 73.85 70.89 72.34 51.43 65.13 60.76 62.87
62.58 74.14 73.20 73.67 54.96 67.79 69.98 68.87
62.66 74.21 73.89 74.05 55.10 68.63 69.88 69.25
68.82 78.13 77.06 77.59 63.30 72.97 75.61 74.27
68.88 79.50 77.06 78.26 63.42 73.03 76.39 74.67
56.28 74.34 54.51 62.90 53.76 72.63 57.44 64.15
61.48 72.03 73.59 72.80 55.69 69.46 68.74 69.10
58.61 73.61 66.20 69.71 52.58 66.34 63.48 64.88
84.90 91.82 86.10 88.87 78.21 85.79 82.56 84.14
84.79 91.74 86.03 88.79 78.16 85.78 82.49 84.10
85.98 91.38 87.04 89.16 76.12 85.79 81.42 83.55
DEFAME 81.23 95.41 79.32 86.63 - - - -
Table 6: Verdict prediction (VP) results on M4FC test set (English-only, imbalanced) by combining intermediate tasks and retrieved evidence (%). ER: RIS evidence retrieval. ✓  and ✓  indicate the inclusion of predicted and ground truth intermediate task outputs and evidence, respectively.
Figure 6: Location verification error examples with Gemini-1.5-Flash on M4FC test. Satellite images are sourced from ESRI World Imagery.
Figure 7: Verdict prediction (VP) F1 scores (%) across different languages. AR: Arabic, NL: Dutch, EN: English, FR: French, DE: German, PT: Portuguese, ES: Spanish, TA: Tamil, TE: Telugu, TR: Turkish.
To further analyze multilingual performance, We disaggregate results by language in Figure 7. Proprietary models again lead overall, generally benefiting from RIS evidence. GPT4o-mini performs weakest on Turkish, while Gemini underperforms on Spanish and Tamil. Telugu is the language on which models perform best, surpassing several higher-resource ones. Yet, open-weight models exhibit smaller or inconsistent gains from evidence, and sometimes perform worse when evidence is added, with Telugu ranking among their weakest languages. These findings suggest that cross-lingual generalization is highly model-dependent, underscoring the need of multilingual evaluation.
Verdict prediction with intermediate tasks
Table 6 presents a detailed evaluation of how intermediate task outputs affects verdict prediction. We consider both predicted and ground truth intermediate outputs, in the imbalanced, English-only setting. We use GPT4o Mini and Qwen2.5-VL-7B, the best performing MLLMs for verdict prediction in the balanced setting (see Figures 4(d)). Figure 1 shows an example of pipeline predictions with GPT4o Mini. Replacing ground truth claims with extracted ones substantially decreases F1, showing that inaccuracies in extracted claims harm downstream verdict prediction. Adding all or some of the other intermediate predicted outputs and the web evidence improves F1. Web evidence has the strongest impact. Surprisingly, using only predicted image contextualization, despite it relying on the same web evidence, performs worse than using the web evidence directly. This suggests that the web evidence is not sufficient to predict accurately most context items during the intermediate image contextualization step. When ground truth intermediate outputs are available, image contextualization yields the largest performance boost, improving F1 by 16.96 percentage points. In that case, adding web evidence provides little to no improvement. The highest F1 is achieved by combining the web evidence and all ground truth intermediate tasks, highlighting the complementarity of the intermediate tasks towards verdict prediction. These results show the promises of combining AFC tasks and open the door for future work on how to best implement these combinations, in terms of accuracy, interpretability, and computational efficiency.
We compare these results with the end to end system DEFAME. DEFAME’s F1 remains lower than the pipeline with ground truth intermediate task outputs. However, it largely outperforms the pipeline with predicted intermediate outputs. This shows that, while intermediate tasks are useful and necessary to produce a fact checking article, SOTA end to end systems like DEFAME still have a clear advantage for verdict prediction.
5 Conclusion
We introduce M4FC, a real-world dataset for multimodal AFC. The images in M4FC have a broad geographic coverage, representing diverse cultures and events. The dataset supports 10 claim languages and spans six tasks, including two new ones, broadening the scope of multimodal AFC. Our baseline results show that M4FC is a challenging dataset for MLLMs and SOTA AFC methods. Furthermore, we showed that intermediate tasks have a high impact on verdict prediction performance.
Limitations
We identify five limitations to M4FC.
First, M4FC does not cover videos and audio. Videos have recently become the most prevalent misinformation modality verified by fact-checkers (Dufour et al., 2024), due to the rise of social media platforms like TikTok. However, obtaining the videos used in real-world multimodal claims is technically and legally challenging. The fact-checking articles usually only include a screenshot, and the videos themselves are often removed from the social media platforms where they appeared. Furthermore, scraping and sharing videos, even for research purposes, is prohibited in many countries.
Second, despite the creation of scraping scripts customized for each fact-checking organization, a lot of manual verification was needed. For example, many image URLs had to be corrected. We also did not cover fact-checking web pages that use JavaScript. This limits the possibilities to scale the dataset. Future work should explore more effective methods for extracting the raw content of fact-checking web pages.
Third, while the images were fact-checked by human experts, the corresponding labels were extracted automatically using GPT4o. As a result, some labels may contain errors. However, evaluations conducted by Prolific annotators indicate that labeling errors introduced by GPT4o are limited.
Fourth, the balanced and multilingual verdict prediction settings suffer from some biases, as discussed in Appendices A. For evaluation, these biases have little to no impact. However, a trained verdict prediction classifier might exploit these biases to achieve artificially high performance. Hence, it is only recommended to use the English-only imbalanced setting for model training.
Fifth, the location verification task contains 195 instances, which is much less than the other tasks. Only a subset of real-world multimodal misinformation cases is suitable for this task, i.e., a ground or aerial view of a location, and a precise location needs to have been identified in the fact-checking article. Hence, scaling the dataset for this task with real-world cases is not a viable option. Instead, future works should consider creating larger synthetic datasets, e.g., by repurposing large news image dataset with location information, such as TARA (Fu et al., 2022).
Ethics statement
Intended use
Research in multimodal AFC tackles the important problem of countering misinformation. By providing a new comprehensive resource, this work enables the development of more performant AFC models. M4FC also contributes to the inclusion of low-resource languages and underrepresented cultural contexts in AFC research. M4FC is intended to be used only for the purpose of multimodal AFC research.
Data access
We collected data from publicly available fact-checking articles. Many images show real-world events, including graphic scenes of violence. Furthermore, the images show real people, some of them identified by their names in the annotations. Throughout the annotation process, we did not contact fact-checkers and claimants. While fact-checkers have often hidden account names and profile pictures from images, they did not do it for all of them. We did not take any additional anonymization steps. For these reasons, we do not directly provide the images. Instead, we provide URLs directing to the images and a script to download them. We also provide a script to download the satellite images and maps needed for the location verification task. We follow the policy of the AVeriTeC dataset (Schlichtkrull et al., 2023): we will remove the annotations for an image-claim pair upon requests of the claimants, the people shown with the image, or the authors of the fact-checking article. The annotations of M4FC are released under a CC BY-SA-4.0 license. The code is released under an Apache 2.0 license.
AI assistants use
AI assistants were used to correct grammar mistakes and typos.
Acknowledgments
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We thank Max Glockner, Shivam Sharma, and Chen Liu for their feedback on a draft of this work.
Appendix A Verdict prediction settings biases
Imbalanced English-only setting
In this default setting, all claims appear as in the fact-checking articles. This setting reflects the real-world distribution of claims verified by fact-checkers, and can be used both for model training and evaluation.
Balanced English-only setting
The balanced English-only setting replaces several false claims with automatically generated true claims based on the true contextual attributes of the image, namely its date, location, event, and depicted people. As a result, the generated true claims typically contain more named entities than real-world claims verified by fact-checkers. In addition, the generated true claims have an average length of 109 characters, whereas real claims in the dataset have an average length of 77 characters. These systematic linguistic differences introduce biases that could be exploited by trained classifiers as superficial heuristics. For this reason, we do not recommend using this setting for model training.
However, we verify that these biases do not affect model evaluation. Table 7 compares verdict prediction performance of two MLLMs under the balanced and imbalanced English-only settings. The resulting differences remain within 3 percentage points. This suggests that, the balanced setting is not less challenging than the imbalanced one as an evaluation setting.
Web evidence Balanced Imbalanced Delta
Qwen2.5VL-7B 69.1 68.8 0.3
72.1 74.3 2.2
GPT4o-mini 74.9 71.9 3.0
77.8 77.6 0.2
Table 7: English-only verdict prediction performance comparison between the balanced and imbalanced settings on M4FC test set, measured with F1 (%). Delta is the absolute difference in F1 score.
Imbalanced multilingual setting
In the multilingual imbalanced setting, most true claims are contained exclusively within two languages, English and Portuguese, while other languages only contain false claims, like Arabic. This is because many non-English speaking fact-checking organizations never write articles about true claims. As a result, a trained model could use the claim’s language as a heuristic for verdict prediction. For this reason, this setting should not be used for model training.
Balanced multilingual setting
In the multilingual balanced setting, both true and false claims are represented for each of the 10 languages. However, the stylistic biases of the synthetic true claims still occurs. For this reason, this setting should not be used for model training.
Appendix B Labeling prompts
Table 8 presents the prompt instructions provided to GPT4o to extract task labels and metadata from the fact-checking articles. Table 9 shows the prompt instructions provided to GPT4o for the post-processing steps.
Please extract the following information from the debunking article and provide the response in the exact format below. Each item should be in one row, and if any information is unavailable, return "None". Do not output anything else.
 Verification strategy: [What verification strategies were used to check the misinformation? (e.g., reverse image search, keyword search, geolocation, deepfake detection, satellite imagery, street view, etc.)]
 Verification tools: [What tools were utilized for verification? (e.g., Google, Yandex, Google Earth, etc.)]
 Claim: [What is the claim accompanying the image?]
 Claimant: [Who made the claim?]
 Claimant’s Intent: [What is the claimant’s intent or agenda behind making the claim?]
 Language of the FC article: [In which language is the fact-checking article written?]
 Language of the claim: [In which language is the claim accompanying the image written?]
 Gold evidence URLs: [List all URLs that link to external webpage evidence.]
 Gold image evidence URLs: [List all URLs linking to image evidence.]
 Claimed date: [According to the claim, when was the image taken?]
 Claimed location: [According to the claim, where was the image taken?]
 Claimed people: [Who are the people shown in the image, according to the claim?]
 Claimed things: [What objects, buildings, plants, or animals are depicted in the image, according to the claim?]
 Claimed event: [What event is being shown in the image, according to the claim?]
 Was the image used before? (Provenance): [Did the fact-checkers retrieve a previous version of the same image? (Answer with Yes, No, or Unknown)]
 Source: [Which person or organization first took or published the image, according to the fact-check? This is different from the claimant who published the claim with the image.]
 Date: [When was the image taken?]
 Location: [Where was the image actually taken?]
 Motivation: [Why was the image taken?]
 People: [Who is shown in the actual image?]
 Things: [What objects, buildings, plants, or animals are shown in the actual image?]
 Event: [What event is depicted in the actual image?]
Here is the debunking article: {Article}
Table 8: GPT4o prompt template to extract task labels and metadata from fact-checking articles.
Post-processing step: Normalize dates to YYYY-MM-DD format
Prompt: Normalize the following dates to YYYY if it is a year, to YYYY-MM if it is a year and month, or YYYY-MM-DD if it is a specific day.
If there is more than one date in the text, repeat the operation for each date, separating them with “;” .
Answer only with the normalized date.
Date: {DATE}
Normalized date:
Post-processing step: Normalizing claims
Prompt: Normalize the following claim about an image by keeping only the actual content.
Remove information about the claimant, unnecessary information such as “the image shows”, and any words indicating the veracity of the claim, such as “viral”, “accurately”, “purported”.
Keep all information that is part of the claim itself.
Answer only with the normalized claim text.
Claim: {CLAIM}
Normalized claim:
Post-processing step: Extracting claim in the original language
Prompt: You are given a fact-checking article debunking a claim, and the translation of this claim in English. Provide the claim as it appears in its original language in the article text. Remove information about the claimant, unnecessary information such as “the image shows”, and any words indicating the veracity of the claim, such as “viral”, “accurately”, “purported” (in the original language).
Answer only with the claim in the original language.
Fact-checking article: {ARTICLE}
Claim: {CLAIM}
Claim in the original language:
Post-processing step: Identify original language
Prompt: In which language is the following claim written. Answer only in one word with the corresponding language.
Claim: {CLAIM}
Language:
Post-processing step: Generate true claims for out-of-context images
Prompt: You are given context information describing an image. Combine this information in one caption of maximum 30 words.
Write the facts only, avoid journalistic style and adjectives, avoid introducing new information.
Date: {DATE}
Location: {LOCATION}
People: {PEOPLE}
Event: {EVENT}
Caption:
Table 9: GPT4o prompt templates to post-process M4FC data.
Appendix C Data collection and labeling details
Figure 8: An example instance provided to the Prolific annotators.
Data scraping
Following the approach of Tonglet et al. (2024), we retrieved all URLs archived on the Wayback Machine555archive.org/help/wayback_api.php under each organization’s web domain. At the time of data collection, all included organizations were current members of the IFCN or under review for membership renewal. We retained URLs containing keywords such as photo, image, or picture, translated into the relevant language. From the raw HTML files, we extracted the article title, publication date, text content, and the URL of the fact-checked image using custom Python scripts. All image URLs were manually verified and corrected when necessary. We manually removed false positives (e.g., non-multimodal claims or fact-checking tutorials), articles containing fact-checking annotations on the image (e.g. an image overlaid with a red cross by the fact-checkers to indicate that it is fake), and duplicate image-claim pairs accross organizations. The final dataset includes 4,982 articles.
GPT4o extraction and post-processing
All GPT4o extracted labels are generated in English and follow a predefined dictionary format. Task labels and metadata that are not available in the article are annotated as “not enough information”. Additional post-processing steps with GPT4o include: (1) normalizing all dates to the YYYY-MM-DD format, (2) rewriting claims to remove any words revealing the verdict, and (3) extracting the claim in its original language in the article and identifying that language. For the multilingual verdict prediction setting, we consider languages represented with at least 40 claims; otherwise, we default to the English translation of the claim. All verdicts are mapped to the following categories: true, false, mixture, unverified. The latter two, which account for only 27 instances, are excluded from verdict prediction. False claims are further categorized into: false (out-of-context), false (manipulated/fake), [^6] Excluding AI-generated images. and false (AI-generated).
Generating synthetic true claims
For each authentic image labeled as false (out-of-context), we generate a synthetic true claim using GPT4o, combining the following context items, if at least one is available: date, location, people, and event. In the balanced verdict setting, we replace the original false claim with the corresponding synthetic true claim in half of the instances where a true claim was successfully generated. The prompt is provided in Appendix B.
Temporal dataset split
The dataset is split temporally into train (60%), dev (10%), and test (30%) sets, using cutoff dates of 2022-09-22, 2023-01-18, and 2024-09-01, respectively, to avoid temporal leakage (Glockner et al., 2022).
Figure 9: The annotators’ interface to validate the labels extracted by GPT4o from fact-checking articles.
Appendix D Data validation - Prolific study
We recruited annotators on the Prolific platform to validate the quality of the data extracted by GPT4o from fact-checking articles. Annotators were paid at the default rate of £9 per hour. We applied the following criteria to select two annotators per study: they need to (1) be fluent speaker of English and their primary language needs to be the one used by the fact-checking organization, (2) have no language-related disorders, (3) have an approval rate of 99% and (4) have already participated in 50 annotation tasks in the past.
D.1 Annotation guidelines
We used the Potato annotation tool (Pei et al., 2022) to design the data validation interface. Each instance consists of the URL of the fact-checking article and the labels extracted by GPT4o, as shown in Figure 8. For each label, annotators have to select either “correct label” or “wrong label”, as shown in Figure 9. Cases where the label is “not enough information” and the annotator selects “wrong label” are automatically mapped to “missing label”, while others are mapped to “incorrect label”.
D.2 Percentage agreement and IAA scores
Table 10 provides the detailed percentages of correct, incorrect, and missing labels, and the IAA scores computed with Randolph’s $\kappa$. We distinguish between labels marked as incorrect or missing. A label is marked as missing when GPT4o outputs “not enough information”, despite relevant information being available. Results are provided per type of label and averaged over all 23 Prolific studies. These results exclude the attention tests. For the percentage of correct, incorrect, and missing labels, we report both the lower bound, where a label is considered correct only if both annotators agree on it, and the upper bound, where an answer is considered correct if at least one annotator agrees on it. IAA is reported both with and without taking into account the cases where the label is “not enough information” (NEI). The scores indicate a clear moderate to high IAA. The highest IAA with NEI is 0.81 for the verdict label. For comparison, the IAA between human annotators for verdict in the AVerImaTeC dataset (Cao et al., 2025) is 0.74.
Label Correct (%) Incorrect (%) Missing (%) Randolph’s $\kappa$
Low Up Low Up Low Up w/o NEI w/ NEI
[…]
Total 75.8 96.6 11.0 0.9 13.3 2.5 - -
Table 10: Detailed results of the Prolific data validation study. Task labels and metadata are highlighted in blue and orange, respectively. Low and Up indicates the lower and upper bound scores, respectively.
Appendix E Additional dataset statistics
Figure 10: Number of M4FC instances with labels for different task combinations.
Figure 11: Cumulative distribution of the verdicts over time. The categories with few instances, unverified, mixture, false (authentic), are not displayed.
Task combination distribution
Figure 10 presents the distribution of task combinations in the dataset. Each row corresponds to a task, and each column represents a unique combination of tasks. The bar chart above indicates the number of instances associated with each combination. The majority of instances are annotated with labels for at least four tasks, while combinations involving only one or two tasks are rare.
Verdict distribution
In the imbalanced setting of verdict prediction, true claims represent only 5.9% of the entire data. This proportion increases to 26.0% in the balanced setting. False claims are dominated by out-of-context misinformation (72%), aligning with prior findings (Dufour et al., 2024). Manipulated and man-made fake images account for 24.0%, while only 3.6% of the false claims involve AI-generated images. Furthermore, AI-generated images mostly appeared from 2023, and are therefore concentrated in the test set. Figure 11 shows the cumulative evolution over time of the verdict labels, with the cutoff dates for the dev and test set indicated by vertical dashed lines.
Claim language distribution
The language distribution under the multilingual setting is as follows: English (63.8%), Arabic (12.4%), Portuguese (5.0%), French (4.6%), Spanish (4.0%), German (3.6%), Telugu (3.0%), Dutch (1.6%), Tamil (1.0%), Turkish (1.0%).
Geographic distribution
Figure 12 compares the geographic distribution of images with location context items in M4FC and two previous datasets with known location labels, VERITE and 5Pils. VERITE is skewed toward Western contexts, while 5Pils is largely concentrated in Asia and Africa. In contrast, M4FC shows the highest geographic diversity, featuring images from a broad range of cultural contexts. Figure 13 further highlights the geographic diversity of M4FC. However, some areas remain underrepresented, such as East Asia, South America, and Oceania. Figure 14 shows the distribution by country of images with known locations. India and the USA are the most represented, which aligns with their large populations and the strong presence of fact-checking organizations.
Figure 12: Distribution by continent of the images with known location of VERITE, 5Pils, and M4FC.
Figure 13: Geographic distribution of images in M4FC with known location.
Figure 14: Top-20 country distribution of images in M4FC with known location. Countries are represented by their ISO 3166-1 alpha-3 codes.
Task # instances # train # dev # test
VCE 223 115 87 21
CIP 124 82 5 37
FD 322 174 31 117
IC 395 215 35 145
LV 12 0 3 9
VP 322 174 31 117
Table 11: Number of instances per task in the human-validated subset.
Figure 15: Distribution by country and by fact-checking organization of the claimed locations of images in M4FC.
Figure 16: Distribution by country and by fact-checking organization of the locations of images in M4FC.
Task: Visual Claim Extraction
Prompt: You are an assistant designed to support fact-checking. Your task is to help identify the primary claim conveyed by the shared image that requires verification. Clearly articulate this claim in a single, concise, and neutral sentence or short paragraph, beginning with “Claim:”
Task: Claimant Intent Prediction
Prompt: You are a helpful assistant supporting fact-checking. Given the image and the related claim: {claim}, your task is to help initialize the prediction of the user’s motivation/intention for sharing this image for further analysis. Clearly articulate it in a single, concise, and neutral sentence or short paragraph, starting with “Motivation: To”
Task: Image Contextualization
Prompt: You are an AI assistant designed to help analyze information. Given the image and related article {evidence}, your task is to identify the following important information: 1. People: Who is shown in the image?
 2. Things: Which animal, plant, building, or object are shown in the image?
 3. Event: Which event is depicted in the image?
 4. Date: When was the image taken?
 5. Location: Where was the image taken?
 6. Motivation: Why was the image taken?
 7. Source: Who is the source of the image? Please provide your answers in the following format in English: People: [Answer],
 Things: [Answer],
 Event: [Answer],
 Date: [Answer],
 Location: [Answer],
 Motivation: [Answer],
 Source: [Answer] If there is no answer, just leave it blank, like an empty string. Do not output anything else. The output must be in English.
Task: Location verification
Prompt: You are a helpful assistant designed to support fact-checking. Your task is to verify whether a candidate location is accurate for an image. You are given the image, and [a map | a satellite | a map and a satellite] view of the candidate location as input. Answer only with true or false.
Task: Verdict prediction Prediction (optionally with text evidence)
Prompt: You are an AI assistant designed to evaluate the veracity of multimodal claims. Given the claim, the related image, and retrieved evidence, your task is to analyze the provided information and predict the more likely veracity: True or False. Please do not refuse or respond with “not enough information.” Your prediction will assist professional fact-checkers in further analysis. Please provide your prediction as “Veracity: True” or “Veracity: False”. In one or more paragraphs, output your reasoning steps. In the separate final line, output your prediction. (Please don’t output anything else except the Veracity Prediction) Claim: {CLAIM} Retrieved Evidences: {EVIDENCE} Let’s think step by step.
Table 12: Prompt templates for the baseline experiments.
Geographic distribution per organization
Figures 15 and 16 show the geographic distribution by country of the claimed and real location of M4FC images, respectively. One map is drawn for each fact-checking organization in M4FC. The maps show that fact-checking organizations tend to focus on images claimed to be associated with their own country. For example, Factly.in and Dfrac.org, both based in India, focus on images with claimed location India. Notable exceptions include Fatabyyano.net, Knack.be, and Dogrula.org. In addition to their own country, many organizations verify images with Ukraine as a claimed location, which reflects the ongoing conflict that started in 2022. For some organizations, the real locations at the country level do not differ significantly from the claimed locations. This is the case for organizations like Factly.in and Indiatoday.in, where all claimed and real locations are within India. This could either mean that location is less frequently misrepresented in out-of-context claims in India, or that the difference in location is at a more fine-grained level, e.g., different cities.
Web evidence statistics
After filtering, we obtain web pages with the Google RIS engine for only 48.62% of the images in M4FC. In total, 9,693 web pages are retrieved, with an average of 4 per image, excluding images with 0 web pages. The web pages are written in 84 different languages. English is the most frequent (65%), followed by Spanish (5%), French (3%), and Arabic (2%).
Appendix F Human-validated subset
We release a separate subset of M4FC, called M4FC-validated, which includes only the labels approved by both Prolific annotators during the data validation phase. This subset contains 434 instances in total. Of these, 235 come from the train set, while 38 and 161 come from the dev and test sets, respectively. The task distribution of this subset is reported in Table 11.
Appendix G G-Eval prompts
Tables 13 and 14 provide the G-Eval prompts for visual claim extraction and claimant intent prediction.
Task: Prompt for accessing visual claim extraction results.
Prompt: You will be given:
1. A ground-truth normalized claim written by a human.
2. An AI-generated normalized claim.
Both claims are produced from the same underlying image content.
Your task is to evaluate how consistent and faithful the AI-generated normalized claim is compared with the ground-truth version.
Definition:
Claim Normalization refers to the process of transforming the information contained in an image into a concise textual claim that captures the core semantic meaning. A good normalized claim should accurately reflect the image’s actual content, entities, relationships, and actions.
Evaluation Criteria:
Consistency (1-5):
5 – Fully faithful to the ground-truth claim; preserves all core semantic elements (entities, relations, actions) with no distortions or additions.
4 – Mostly faithful; minor omissions or rephrasing but no incorrect or contradictory content.
3 – Partially faithful; preserves some main elements but misses key relationships or adds ambiguous wording.
2 – Low faithfulness; multiple incorrect or missing semantic elements; contains hallucinations or distorts relationships.
1 – Poor; contradicts the main meaning of the ground-truth claim or fails to represent the image content accurately.
Evaluation Steps:
1. Read the ground-truth normalized claim and identify its core semantic units:
- entities (who/what)
- attributes (descriptions)
- actions/events
- relationships
- spatial or contextual cues
2. Read the AI-generated normalized claim.
3. Compare the semantic overlap and accuracy:
- Does the AI version preserve the core entities?
- Does it distort actions or relationships?
- Does it add hallucinated content not in the ground-truth?
- Does it omit essential elements?
4. Assign a Consistency score (1-5) strictly based on the Evaluation Criteria.
Return STRICT JSON with keys:
{"verdict": one of [consistent, partially_consistent, inconsistent], "score": integer 1-5, "explanation": brief reason}. No extra text.
Evaluation Form (scores ONLY):
- Consistency:
Ground Truth Normalized Claim:
<ground-truth-claim-here>
AI-Generated Normalized Claim:
<ai-generated-claim-here>
Respond with JSON only.
Table 13: G-Eval prompt for VCE task.
Task: Prompt for accessing claimant intent prediction results.
Prompt: You will be given:
1. A ground-truth motivation written by a human.
2. An AI-generated motivation.
Your task is to evaluate how well the AI-generated motivation is consistent and faithful to the meaning and reasoning of the ground-truth version.
Evaluation Criteria:
Consistency (1-5):
5 – Fully faithful, no contradictions, and captures all essential reasoning points.
4 – Mostly faithful; minor omissions but no distortions.
3 – Partially faithful; captures some main points but misses or alters key logic.
2 – Low faithfulness; contains distortions, irrelevant reasoning, or major omissions.
1 – Poor; contradicts or misrepresents the ground-truth motivation.
Evaluation Steps:
1. Read the ground-truth motivation and identify its key reasoning points.
2. Read the AI-generated motivation.
3. Compare meaning, reasoning steps, and any contradictions or hallucinations.
4. Assign a Consistency score from 1 to 5 based on the Evaluation Criteria.
Return STRICT JSON with keys:
{"verdict": one of [consistent, partially_consistent, inconsistent], "score": integer 1-5, "explanation": brief reason}. No extra text.
Evaluation Form (scores ONLY):
- Consistency:
Ground Truth Motivation:
<ground-truth-text-here>
AI-Generated Motivation:
<ai-generated-text-here>
Respond with JSON only.
Table 14: G-Eval prompt for CIP task.
Appendix H Prompt templates for the M4FC tasks
Table 12 reports the prompt templates used for the baseline experiments.
Appendix I Implementation details
MLLMs
All models are evaluated in one zero-shot run on one A100 GPU with temperature set to 0. We use the HuggingFace (Wolf et al., 2020) implementation of the open-weight models. GPT4o-mini and Gemini-1.5-Flash are accessed via the OpenAI API and Google Cloud API, respectively.
Fake image detectors fine-tuning
We fine-tune all models with the Adam optimizer and a learning rate of 2e52e-5 for 10 epochs.
DEFAME
We employ the Google Custom Search API for both text and image retrieval. We allow up to three iterations, with each retrieval step returning the top-3 evidence items. The maximum output length of a single LLM generation is limited to 2,000 tokens (max_tokens), while the overall fact-checking process, including the final report and all intermediate steps, is capped at 64,000 tokens (max_result_len).
COVE
We reproduce all steps of the COVE pipeline, except the collection of evidence via direct search, which has limited impact on real-world misinformation detection (Tonglet et al., 2025).
Rouge-L BERTScore METEOR
InternVL2.5-8B 25.70 51.43 18.04
Qwen2.5-VL-7B 28.44 58.21 20.09
Llama-3.2-11B 19.82 48.73 12.01
GPT4o-mini 26.94 56.2 18.70
Gemini-1.5-Flash 32.78 64.7 23.67
Table 15: Visual claim extraction results (%).
Rouge-L BERTScore METEOR
InternVL2.5-8B 15.30 32.20 8.04
Qwen2.5-VL-7B 15.35 34.81 7.93
Llama-3.2-11B 11.76 32.04 5.43
GPT4o-mini 11.14 30.23 5.47
Gemini-1.5-Flash 17.13 45.72 8.70
Table 16: Claimant intent prediction results (%).
Model Acc F1 Pre Rec
Deep-Fake Detector [Uncaptioned image] 58.05 70.31 64.05 77.92
Deep-Fake Detector [Uncaptioned image] 69.49 78.88 70.58 89.38
AI-Image Detector [Uncaptioned image] 36.74 6.41 56.14 3.40
AI-Image Detector [Uncaptioned image] 68.34 78.49 69.21 90.66
RealFake Detector [Uncaptioned image] 55.48 60.17 70.00 52.76
RealFake Detector [Uncaptioned image] 67.87 79.64 76.16 83.47
Table 17: Fake image detection results (%). [Uncaptioned image] and [Uncaptioned image] indicate predictions of the base pre-trained model and of the model fine-tuned on M4FC train set, respectively.
RIS evidence
For each image, we retrieve up to 20 web pages with the Google RIS engine. We extract their textual content using Trafilatura (Barbaresi, 2021).777trafilatura.readthedocs.io To prevent evidence leakage, we exclude any web pages published by a curated list of fact-checking organizations, as well as those published after the date of the fact-checking article from which the claim was sourced. For tasks that require text evidence, we treat each web page retrieved with RIS as a single evidence. All 20 text evidence are provided in the prompt, sorted by cosine similarity with the image using CLIP embeddings (Radford et al., 2021).
Appendix J Additional Results
Tables 15,16, 17, and 18 show the results tables on M4FC test corresponding to Figures 4 and 7.
Language GPT4o-mini Gemini-1.5-Flash Qwen2.5-VL-7B InternVL2.5-8B LLaMA-3.2-11B
w/o evi w/ evi w/o evi w/ evi w/o evi w/ evi w/o evi w/ evi w/o evi w/ evi
Arabic 80.11 81.09 63.11 70.13 68.36 73.06 50.52 52.41 31.67 42.57
Dutch 72.56 82.43 65.73 62.95 73.48 74.68 55.62 57.37 40.21 44.54
English 76.12 78.14 68.44 76.34 71.52 74.58 64.12 65.19 60.68 58.32
French 73.25 80.04 68.28 74.28 73.89 81.25 61.78 67.67 43.26 43.29
German 69.07 68.27 64.34 69.52 74.47 81.36 55.93 64.40 47.83 44.36
Portuguese 66.73 73.42 67.12 79.10 72.71 79.30 59.65 63.84 47.23 43.24
Spanish 73.86 78.87 60.26 64.17 63.40 56.55 62.56 44.85 47.69 12.31
Tamil 78.37 80.22 57.55 62.32 72.92 71.65 52.20 60.32 57.82 37.40
Telugu 90.32 90.91 72.48 84.23 84.05 86.73 52.60 41.21 31.59 32.78
Turkish 52.16 71.49 67.28 72.62 76.76 65.37 62.10 35.98 30.20 28.08
Table 18: Verdict prediction (multilingual, balanced) F1 per language, with (w/) and without (w/o) evidence (%).
Zoom [Uncaptioned image] [Uncaptioned image] F1
10 68.85
10 66.59
10 64.14
15 74.43
15 67.52
15 65.62
17 73.63
17 70.55
17 69.75
Table 19: Location verification F1 at different zoom level on M4FC test set with Gemini-1.5-Flash (%).
Figure 17: Illustration of location verification with satellite views at different zoom level. Satellite images are sourced from ESRI World Imagery.
Appendix K Location verification at different zoom levels
As shown in the second error example of Figure 6, better results could be obtained at different zoom levels, as illustrated in Figure 17. However, the appropriate zoom level may vary a lot. Future work should consider directly incorporating the desired zoom level for the aerial views as a parameter of their modeling approach. We provide preliminary results in Table 19, showing the F1 of Gemini-1.5-Flash on M4FC test set at three different zoom levels: 10, 15, and 17. The overall best performance is obtained with a zoom level of 15.